### v6

@@ Appendix A Some details in the Proofs of Well-posedness and Mean-field limit
Proof B.1.
@@
where the first inequality comes from Assumption 4.2 (A2) and the second inequality comes from Assumption 4.2 (C1). Then by taking $\tau_{2}(x)$ to be $\tau_{1}^{\beta}(x)$ will finish the proof of the existence of $\tau_{2}$.
@@
where in the first inequality, we used Assumption 4.2 (A2) and in the third inequality, we used Assumption 4.2 (C1). Then one has
@@
Therefore, selecting $\tau_{3}(x)$ as a scalar multiple of $2x^{\beta}+\tau_{1}^{\beta}(x)$ will suffice. One can apply the same method to prove the existence of $\tau_{4}(x)$.
Appendix C Explanations for expanding the test function space
@@
$d\phi(\bar{V}_{t})=$ $\nabla\phi\left(\bar{V}_{t}\right)\cdot\left(\left(-\lambda\left(\bar{V}_{t}-v_{\alpha}\left(\rho_{t}\right)\right)-\dfrac{1}{\epsilon}\nabla G\left(\bar{V}_{t}\right)\right)\,dt\right)$
$+\dfrac{1}{2}\sigma^{2}\sum_{k=1}^{d}\partial_{kk}\phi\left(\bar{V}_{t}\right)\left(\bar{V}_{t}-v_{\alpha}\left(\rho_{t}\right)\right)_{k}^{2}\,dt+\sigma\nabla\phi\left(\bar{V}_{t}\right)^{T}\text{diag}\left(\bar{V}_{t}-v_{\alpha}\left(\rho_{t}\right)\right)\,dB_{t}.$
Note that $\mathbb{E}\int_{0}^{t}\sigma\nabla\phi\left(\bar{V}_{t}\right)^{T}\text{diag}\left(\bar{V}_{t}-v_{\alpha}\left(\rho_{t}\right)\right)\,dB_{t}=0$ by applying Theorem 3.2.1 (iii) in [34] due to the facts that $\phi\in\mathcal{C}_{*}^{2}\left(\mathbb{R}^{d}\right)$ and $\rho_{t}\in\mathcal{C}\left([0,T],\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)\right)$. Taking the expectation and applying Fubini’s theorem gives
$\dfrac{d}{dt}\mathbb{E}\phi\left(\bar{V}_{t}\right)=$ $-\lambda\mathbb{E}\nabla\phi\left(\bar{V}_{t}\right)\cdot\left(-\lambda\left(\bar{V}_{t}-v_{\alpha}\left(\rho_{t}\right)\right)-\dfrac{1}{\epsilon}\nabla G\left(\bar{V}_{t}\right)\right)$
$+\dfrac{1}{2}\sigma^{2}\mathbb{E}\sum_{k=1}^{d}\partial_{kk}\phi\left(\bar{V}_{t}\right)\left(\bar{V}_{t}-v_{\alpha}\left(\rho_{t}\right)\right)_{k}^{2},$
@@
Appendix D Proof of Lemma 4.5
Proof D.1.
@@
(62) $\dfrac{d}{dt}\mathcal{V}\left(\rho_{t}\right)=-\lambda\int\Bigl{\langle}v-v_{\alpha}\left(\rho_{t}\right),v-v^{*}\Bigr{\rangle}\,d\rho_{t}-\dfrac{1}{\epsilon}\int\Bigl{\langle}\nabla G,v-v^{*}\Bigr{\rangle}\,d\rho_{t}+\dfrac{\sigma^{2}}{2}\int\|v-v(\rho_{t})\|^{2}\,d\rho_{t}.$
@@
$=-\lambda\int\Bigl{\langle}v-v^{*},v-v^{*}\Bigr{\rangle}\,d\rho_{t}(v)+\lambda\int\Bigl{\langle}v-v^{*},v_{\alpha}\left(\rho_{t}\right)-v^{*}\Bigr{\rangle}\,d\rho_{t}(v)$
@@
(63) $-\lambda\int\Bigl{\langle}v-v_{\alpha}\left(\rho_{t}\right),v-v^{*}\Bigr{\rangle}\,d\rho_{t}(v)\leq-2\lambda\mathcal{V}\left(\rho_{t}\right)+\lambda\|\int(v-v^{*})\,d\rho_{t}(v)\|_{2}\cdot\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$
@@
$=$ $\dfrac{\sigma^{2}}{2}\left(\int\|v-v^{*}\|_{2}^{2}\,d\rho_{t}(v)-2\Bigl{\langle}\int(v-v^{*})\,d\rho_{t}(v),v_{\alpha}\left(\rho_{t}\right)-v^{*}\Bigr{\rangle}+\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}^{2}\right)$
$\leq$ $\sigma^{2}\left(\mathcal{V}\left(\rho_{t}\right)+\int\|v-v^{*}\|_{2}\,d\rho_{t}(v)\cdot\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}+\dfrac{1}{2}\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}^{2}\right)$
$\leq$ $\sigma^{2}\left(\mathcal{V}\left(\rho_{t}\right)+\sqrt{2\mathcal{V}\left(\rho_{t}\right)}\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}+\dfrac{1}{2}\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}^{2}\right),$
where in the first inequality, we use Cauchy’s inequality and Minkowski’s inequality and in the second inequality, we use Cauchy’s inequality again. Plugging (63) and (64) back into (62) finishes the proof.
Appendix E Lemmas used in Laplace’s principle
E.1 Lemma E.1 and proof
Lemma E.1.
@@ Appendix E Lemmas used in Laplace’s principle
$\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}\right)^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v).$
Here, $\mathcal{E}^{0}_{r}$, $\mathcal{E}_{\infty}$ and $\tau_{3}$ are quantities defined in Assumption 4.2 (C) and Lemma 4.6.
Proof E.2.
@@
By Assumption 4.2 (C1), $\partial B^{\infty}(v^{*},r)\cap\{G=0\}\neq\emptyset$, which leads to
@@
where the first inequality comes from Assumption 4.2 (C2), the second inequality is due to $v\in B^{\infty}(v^{*},\tilde{r})^{c}$, the third inequality is because of the definition of $\tilde{r}$ and the last equality is because we assumed $\mathcal{E}(v^{*})=0$. This completes the proof of the second one.
@@
(65) $\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})}\dfrac{\big{\|}v-v^{*}\big{\|}_{\infty}}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\tilde{r}\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})}\dfrac{1}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\tilde{r}$.
@@
$\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\rho_{t})}$ $=\int e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\geq\int_{B^{\infty}(v^{*},r)}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\geq\int_{B^{\infty}(v^{*},r)}e^{-\alpha\mathcal{E}_{r}}\,d\rho_{t}(v)$
@@
$\int_{\{G=0\}}\dfrac{\big{\|}v-v^{*}\big{\|}_{\infty}}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\dfrac{(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\big{\|}v-v^{*}\big{\|}_{\infty}\,d\rho_{t}(v).$
@@
$\int_{\{G=0\}}\dfrac{\big{\|}v-v^{*}\big{\|}_{2}}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\dfrac{\sqrt{d}(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\big{\|}v-v^{*}\big{\|}_{2}\,d\rho_{t}(v).$
This completes the proof.
E.2 Lemma E.3 and proof
Lemma E.3.
@@
$\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq$ $\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}$
$+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{t}(v)$
$+\sqrt{d}\tau_{1}(u).$
Here, $v_{G(v)}=\arg\min_{v^{\prime}\in\{G(v^{\prime})=G(v)\}}\mathcal{E}(v^{\prime})$, $\underline{\mathcal{E}_{\tilde{u}}}$ and $\tau_{1}$ are defined in Assumption 4.2 (C1), $\mathcal{E}_{r}^{\tilde{u}}$, $\tau_{3}$ and $\tau_{4}$ are quantities defined in Lemma 4.6.
Proof E.4.
@@
$\int_{\{G\in(0,u)\}}\dfrac{\big{\|}v-v^{*}\big{\|}_{\infty}}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$ $=\int_{\{G\in(0,u)\}}\dfrac{\big{\|}v-v^{*}\big{\|}_{\infty}}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\rho_{t})}\big{\|}\nabla G\big{\|}_{2}}e^{-\alpha\mathcal{E}(v)}\big{\|}\nabla G\big{\|}_{2}\rho_{t}\,dv$

### v8

@@ Appendix A Some details in the Proofs of Well-posedness and Mean-field limit
Proof.
We verify Assumption 2 (B) one by one.
Verification of Assumption 2 (B1)
We first prove when $K$ is convex, the distance function $g$ is convex. Given $\lambda_{1},\lambda_{2}>0$ with $\lambda_{1}+\lambda_{2}=1$ and $v_{1},v_{2}\in\mathbb{R}^{d}$. By the definition of distance function, for any $\epsilon>0$, there exist $u_{1}$ and $u_{2}$ in $K$ such that,
$\|v_{1}-u_{1}\|_{2}\leq g(v_{1})+\epsilon,$
$\|v_{2}-u_{2}\|_{2}\leq g(v_{2})+\epsilon.$
Thus one has
$g(\lambda_{1}v_{1}+\lambda_{2}v_{2})$ $\leq\|\lambda_{1}v_{1}+\lambda_{2}v_{2}-\lambda_{1}u_{1}-\lambda_{2}u_{2}\|_{2}$
$\leq\lambda_{1}\|v_{1}-u_{1}\|_{2}+\lambda_{2}\|v_{2}-u_{2}\|_{2}$
$\leq\lambda_{1}g(v_{1})+\lambda_{2}g(v_{2})+\epsilon.$
The above inequality holds for all $\epsilon>0$. This gives the convexity of $g$. Also, one knows that $x^{2}$ is a convex increasing function when $x\geq 0$. As a result, the composition $G=g^{2}$ is convex, thereby satisfying Assumption 2 (B1).
Verification of Assumption 2 (B2)
Since we are dealing with a compact convex set $K\subset\mathbb{R}^{d}$, we can conveniently work with the Moreau envelope. Let $P_{K}$ denote the metric projection onto $K$. Then
$G(v)=\operatorname{dist}(v,K)^{2}=\|v-P_{K}(v)\|^{2},\qquad v\in\mathbb{R}^{d}.$
Let $\iota_{K}$ be the indicator function of $K$, that is,
$\iota_{K}(v)=\begin{cases}0,&v\in K,\\ +\infty,&v\notin K.\end{cases}$
The Moreau envelope of $\iota_{K}$ is defined by
$M_{\iota_{K}}(u)=\inf_{v\in\mathbb{R}^{d}}\biggl\{\iota_{K}(v)+\frac{1}{2}\,\|v-u\|^{2}\biggr\}.$
Since $\iota_{K}(v)=0$ for $v\in K$ and $\infty$ otherwise, we obtain
$M_{\iota_{K}}(u)=\inf_{v\in K}\frac{1}{2}\,\|v-u\|^{2}=\frac{1}{2}\,\|u-P_{K}(u)\|^{2}=\frac{1}{2}G(u),$
so $M_{\iota_{K}}=G/2$.
We now use [6, Proposition 12.30], which yields
$\nabla M_{\iota_{K}}(u)=u-\mathrm{prox}_{\iota_{K}}(u),$
where $\mathrm{prox}_{\iota_{K}}$ is the proximal operator of $\iota_{K}$. In the present case one has $\mathrm{prox}_{\iota_{K}}=P_{K}$, and therefore
$\nabla G(v)=2\nabla M_{\iota_{K}}(v)=2\bigl(v-P_{K}(v)\bigr),$
so that
$\|\nabla G(v)\|^{2}=4\|\nabla M_{\iota_{K}}(v)\|^{2}=4\|v-P_{K}(v)\|^{2}=4G(v),$
for all $v\in\mathbb{R}^{d}$. Hence the inequality in Assumption 2 (B2) is satisfied with equality.
It remains to prove that $G\in\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$. In other words, we need to show that $|\partial_{k}G(x)|\leq C(1+|x_{k}|)$ and $\sup_{x\in\mathbb{R}^{d}}|\partial_{kk}G(x)|<\infty$ for all $k=1,\dots,d$.
The first bound follows immediately from the expression $\nabla G(v)=2\bigl(v-P_{K}(v)\bigr)$:
$|\partial_{k}G(v)|=2\,|v_{k}-P_{K}(v)_{k}|\leq 2\bigl(|v_{k}|+\sup_{u\in K}|u_{k}|\bigr)\leq C(1+|v_{k}|),$
where $\sup_{u\in K}|u_{k}|<\infty$ because $K$ is compact, and $C>0$ is a constant independent of $v$ and of the dimension.
For the second bound, we recall that the gradient of a Moreau envelope is always globally Lipschitz; in particular, $\nabla M_{\iota_{K}}$ is 11–Lipschitz, hence $\nabla G=2\nabla M_{\iota_{K}}$ is globally Lipschitz. By Rademacher’s theorem, $G$ is twice differentiable almost everywhere, and the Hessian $\nabla^{2}G(x)$ exists for almost every $x$ with operator norm bounded by the Lipschitz constant of $\nabla G$. Thus $\|\nabla^{2}G(x)\|_{\mathrm{op}}$ is essentially bounded, and the diagonal entries $\partial_{kk}G(x)$ (which are nonnegative due to the convexity of $G$) are uniformly bounded as well. This yields
$\sup_{x\in\mathbb{R}^{d}}|\partial_{kk}G(x)|<\infty,\qquad k=1,\dots,d.$
Hence $G\in\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$, and Assumption 2 (B2) is verified. Moreover, all the constants involved can be chosen independently of the ambient dimension.
Verification of Assumption 2 (B3)
We first recall the relevant notions from convex geometry.
Definition B.1 (Intrinsic volumes, [34, Definition 1.1]).
Let $S\subset\mathbb{R}^{d}$ be a nonempty compact convex set. For each index $j=0,1,\dots,d$, let $P_{j}\in\mathbb{R}^{d\times d}$ be the orthogonal projector onto a fixed jj–dimensional subspace of $\mathbb{R}^{d}$, and let $Q\in\mathbb{R}^{d\times d}$ be a random rotation matrix distributed according to the Haar measure on the special orthogonal group $\mathrm{SO}(d)$. The intrinsic volumes of $S$ are defined by
$V_{j}(S):=\binom{d}{j}\frac{\kappa_{d}}{\kappa_{j}\kappa_{d-j}}\,\mathbb{E}_{Q}\bigl[\operatorname{Vol}_{j}(P_{j}QS)\bigr],\qquad j=0,\dots,d,$
where $\kappa_{d}$ denotes the volume of the dd–dimensional Euclidean unit ball and $\operatorname{Vol}_{j}(\cdot)$ is the jj–dimensional Lebesgue measure computed relative to the affine hull. In particular, $V_{d}(S)$ coincides with the dd–dimensional Lebesgue measure of $S$.
Definition B.2 (Parallel sets).
Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set and let $\operatorname{dist}(\cdot,K)$ denote the Euclidean distance to $K$. For $r\geq 0$ the parallel set of $K$ at distance $r$ is defined by
$K_{\leq r}:=\{v\in\mathbb{R}^{d}:\operatorname{dist}(v,K)\leq r\}.$
We will use the following standard results.
Proposition B.3 (Steiner formula, [34, Fact 2.1]).
Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set. For every $r\geq 0$ one has
$V_{d}(K_{\leq r})=\sum_{j=0}^{d}\kappa_{d-j}V_{j}(K)\,r^{d-j}.$
Proposition B.4 ([42, Corollary 2.5]).
Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set. For all $r>0$, except possibly for a countable set of values of $r$, one has
$\frac{d}{dr}V_{d}(K_{\leq r})=\mathcal{H}^{d-1}(\partial K_{\leq r}),$
where $\mathcal{H}^{d-1}(\cdot)$ denotes the (d1)(d-1)–dimensional Hausdorff measure.
We now prove Assumption 2 (B3) holds. Recall that $G=g^{2}$ with $g(v)=\operatorname{dist}(v,K)$, and from the verification of (B2) we have
$\|\nabla G(v)\|=\sqrt{4G(v)}=2g(v),\qquad v\in\mathbb{R}^{d}.$
At points where $g$ is differentiable and $g(v)>0$,
$\nabla G(v)=2g(v)\,\nabla g(v),$
so
$2g(v)=\|\nabla G(v)\|=\|2g(v)\,\nabla g(v)\|=2g(v)\,\|\nabla g(v)\|,$
and hence $\|\nabla g(v)\|=1$ at all such points.
By the coarea formula applied to the Lipschitz function $g$ we obtain
$\int_{\{G\in(0,1)\}}\frac{1}{\|\nabla G(v)\|}\,dv$ $=\int_{\{G\in(0,1)\}}\frac{1}{2g(v)}\,\|\nabla g(v)\|\,dv$
$=\int_{0}^{1}\int_{\{g(v)=t\}}\frac{1}{2g(v)}\,d\mathcal{H}^{d-1}(v)\,dt$
$=\int_{0}^{1}\frac{1}{2t}\,\mathcal{H}^{d-1}\bigl(\{v\in\mathbb{R}^{d}:g(v)=t\}\bigr)\,dt$
$=\int_{0}^{1}\frac{1}{2t}\,\mathcal{H}^{d-1}(\partial K_{\leq t})\,dt,$
where in the last step we used $\{v:g(v)=t\}=\partial K_{\leq t}$.
From Propositions B.3 and B.4 we have, for all $t>0$ outside a countable set,
$\mathcal{H}^{d-1}(\partial K_{\leq t})$ $=\frac{d}{dt}V_{d}(K_{\leq t})$
$=\sum_{j=0}^{d-1}(d-j)\,\kappa_{d-j}\,V_{j}(K)\,t^{d-j-1}.$
Therefore,
$\int_{\{G\in(0,1)\}}\frac{1}{\|\nabla G(v)\|}\,dv=\frac{1}{2}\sum_{j=0}^{d-1}(d-j)\,\kappa_{d-j}\,V_{j}(K)\int_{0}^{1}t^{d-j-2}\,dt.$
For $j\leq d-2$ the integral $\int_{0}^{1}t^{d-j-2}\,dt$ is finite, whereas for j=d1j=d-1 the integrand behaves like $t^{-1}$ near 0 and the integral diverges. Hence finiteness of the above expression requires
$V_{d-1}(K)=0.$
A sufficient condition for this is that $K$ is contained in an affine subspace of dimension less than or equal to d2d-2, in which case the (d1)(d-1)–dimensional intrinsic volume $V_{d-1}(K)$ vanishes. Under this condition, Assumption 2 (B3) holds.
Appendix C Proof of Lemma 4.4
Proof.
@@
where the first inequality comes from Assumption 2 (A2) and the second inequality comes from Assumption 2 (C1). Then by taking $\tau_{2}(x)$ to be $\tau_{1}^{\beta}(x)$ will finish the proof of the existence of $\tau_{2}$.
@@
where in the first inequality, we used Assumption 2 (A2) and in the third inequality, we used Assumption 2 (C1). Then one has
@@
Therefore, selecting $\tau_{3}(x)$ as a scalar multiple of $2x^{\beta}+\tau_{1}^{\beta}(x)$ will suffice. One can apply the same method to prove the existence of $\tau_{4}(x)$. ∎
Appendix D Explanations for expanding the test function space
@@
$d\phi(\bar{V}_{t})=$ $\nabla\phi\left(\bar{V}_{t}\right)\cdot\left(\left(-\lambda\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)-\dfrac{1}{\epsilon}\nabla G\left(\bar{V}_{t}\right)\right)\,dt\right)$
$+\dfrac{1}{2}\sigma^{2}\sum_{k=1}^{d}\partial_{kk}\phi\left(\bar{V}_{t}\right)\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\,dt+\sigma\nabla\phi\left(\bar{V}_{t}\right)^{T}\text{diag}\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)\,dB_{t}.$
Note that $\mathbb{E}\int_{0}^{t}\sigma\nabla\phi\left(\bar{V}_{t}\right)^{T}\text{diag}\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)\,dB_{t}=0$ by applying Theorem 3.2.1 (iii) in [39] due to the facts that $\phi\in\mathcal{C}_{*}^{2}\left(\mathbb{R}^{d}\right)$ and $\rho_{t}\in\mathcal{C}\left([0,T],\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)\right)$. Taking the expectation and applying Fubini’s theorem gives
$\dfrac{d}{dt}\mathbb{E}\phi\left(\bar{V}_{t}\right)=$ $-\lambda\mathbb{E}\nabla\phi\left(\bar{V}_{t}\right)\cdot\left(-\lambda\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)-\dfrac{1}{\epsilon}\nabla G\left(\bar{V}_{t}\right)\right)$
$+\dfrac{1}{2}\sigma^{2}\mathbb{E}\sum_{k=1}^{d}\partial_{kk}\phi\left(\bar{V}_{t}\right)\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2},$
@@
Appendix E Proof of Lemma 4.3
Proof.
@@
(66) $\dfrac{d}{dt}\mathcal{V}\left(\rho_{t}\right)=$ $-\lambda\int\Bigl\langle v-v_{\alpha}\left(\rho_{t}\right),v-v^{*}\Bigr\rangle\,d\rho_{t}$
$-\dfrac{1}{\epsilon}\int\Bigl\langle\nabla G,v-v^{*}\Bigr\rangle\,d\rho_{t}+\dfrac{\sigma^{2}}{2}\int\|v-v(\rho_{t})\|^{2}\,d\rho_{t}.$
@@
$=-\lambda\int\Bigl\langle v-v^{*},v-v^{*}\Bigr\rangle\,d\rho_{t}(v)+\lambda\int\Bigl\langle v-v^{*},v_{\alpha}\big(\rho_{t}\big)-v^{*}\Bigr\rangle\,d\rho_{t}(v)$
@@
(67) $-\lambda\int\Bigl\langle v-v_{\alpha}\left(\rho_{t}\right),v-v^{*}\Bigr\rangle\,d\rho_{t}(v)$
$\leq-2\lambda\mathcal{V}\left(\rho_{t}\right)+\lambda\|\int(v-v^{*})\,d\rho_{t}(v)\|_{2}\cdot\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$
@@
$=$ $\dfrac{\sigma^{2}}{2}\left(\int\|v-v^{*}\|_{2}^{2}\,d\rho_{t}(v)-2\Bigl\langle\int(v-v^{*})\,d\rho_{t}(v),v_{\alpha}\left(\rho_{t}\right)-v^{*}\Bigr\rangle+\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}\right)$
$\leq$ $\sigma^{2}\left(\mathcal{V}\big(\rho_{t}\big)+\int\|v-v^{*}\|_{2}\,d\rho_{t}(v)\cdot\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}+\dfrac{1}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}\right)$
$\leq$ $\sigma^{2}\left(\mathcal{V}\big(\rho_{t}\big)+\sqrt{2\mathcal{V}\big(\rho_{t}\big)}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}+\dfrac{1}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}\right),$
where in the first inequality, we use Cauchy’s inequality and Minkowski’s inequality and in the second inequality, we use Cauchy’s inequality again. Plugging (67) and (68) back into (66) finishes the proof. ∎
Appendix F Lemmas used in Laplace’s principle
In this section, we present the two lemmas (Lemma F.1 and Lemma F.2) used in the proof of Lemma 4.5.
F.1. Lemma F.1 and proof
Lemma F.1.
@@ Appendix F Lemmas used in Laplace’s principle
$\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}\right)^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v).$
Here, $\mathcal{E}^{0}_{r}$, $\mathcal{E}_{\infty}$ and $\tau_{3}$ are quantities defined in Assumption 2 (C) and Lemma 4.4.
Proof.
@@
By Assumption 2 (C1), $\partial B^{\infty}(v^{*},r)\cap\{G=0\}\neq\emptyset$, which leads to
@@
where the first inequality comes from Assumption 2 (C2), the second inequality is due to $v\in B^{\infty}(v^{*},\tilde{r})^{c}$, the third inequality is because of the definition of $\tilde{r}$ and the last equality is because we assumed $\mathcal{E}(v^{*})=0$. This completes the proof of the second one.
@@
(69) $\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\tilde{r}\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\tilde{r}$.
@@
$\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}$ $=\int e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\geq\int_{B^{\infty}(v^{*},r)}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\geq\int_{B^{\infty}(v^{*},r)}e^{-\alpha\mathcal{E}_{r}}\,d\rho_{t}(v)$
@@
$\int_{\{G=0\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\big\|v-v^{*}\big\|_{\infty}\,d\rho_{t}(v).$
@@
$\int_{\{G=0\}}\dfrac{\big\|v-v^{*}\big\|_{2}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\big\|v-v^{*}\big\|_{2}\,d\rho_{t}(v).$
This completes the proof. ∎
F.2. Lemma F.2 and proof
Lemma F.2.
@@
$\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\quad+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{t}(v)$
$\quad+\sqrt{d}\tau_{1}(u).$
Here, $v_{G(v)}=\arg\min_{v^{\prime}\in\{G(v^{\prime})=G(v)\}}\mathcal{E}(v^{\prime})$, $\underline{\mathcal{E}_{\tilde{u}}}$ and $\tau_{1}$ are defined in Assumption 2 (C1), $\mathcal{E}_{r}^{\tilde{u}}$, $\tau_{3}$ and $\tau_{4}$ are quantities defined in Lemma 4.4.
Proof.
@@
$=\int_{\{G\in(0,u)\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\big\|\nabla G\big\|_{2}\rho_{t}\,dv$